THE GROUND STATE TORSION-ROTATION SPECTRUM OF PROPARGYL ALCOHOL (HCCCH_{2}OH)

\underline{J. C. PEARSON}, Jet Propulsion Laboratory, California Institute of Technology, Mail Stop 183-301, 4800 Oak Grove Dr., Pasadena, CA 91109.

The ground state torsion-rotation spectrum of the asymmetric internal hydroxyl rotation in propargyl alcohol (HCCCH_{2}OH) has been investigated in the 84 to 640 GHz range. The ground state has been confirmed to consists of two torsional sub-states of the gauche conformer^{a}, a symmetric gauche state (gauche+) and an antisymmetric gauche state (gauche-). No evidence of a low lying third state or trans conformer has been observed. The more stable gauche+ state is 652389.5 MHz below the gauche- state. Due to their close proximity, the two gauche states interact strongly through a series of a- and b-type Coriolis interactions. Strong a- and weak b-dipole rotational transitions are observed within each sub-state while strong c-dipole torsional transitions are observed between the two sub-states. Over 2300 transitions covering a range of J and K_{a} values to 80 and 34, respectively, have been fit to experimental accuracy using a fixed-frame-axis method (FFAM) Hamiltonian. The rotation, distortion and interaction constants have been determined. The 652 GHz gauche+-gauche- energy difference, strong gauche+ to gauche- transitions and the constants derived from the analysis should enable rapid astronomical detection.

aEizi Hirota, J. Mol. Spectrosc. \underline{26}, 335 (1968).